OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

OpenStudy (anonymous):

-x+2y=9 -x=9-2y x=2y-9 or x=-9+2y, so the first option :D

OpenStudy (anonymous):

thank you